# PRESERVE: Prefetching Model Weights and KV-Cache in Distributed LLM Serving

### v1

@@ 5 Design Space Exploration
Table 2: Model parameters used in design space exploration.
@@
Table 2 summarizes the values of the model parameters used in our design space exploration. The silicon area for cores and L2 SRAM cache are taken from Lin et al. [24], whose design is fabricated in a 7nm technology node. We assumed that the accelerator has a theoretical peak throughput of 800 TeraOps/s, which is similar to existing inference accelerators [12, 21] and it is equipped with 4x HBM2e memory, which provides 64 GB of capacity and 1840 GB/s bandwidth in total [33].
For the design space exploration, we selected a number of open-source, state-of-the-art LLMs, namely Llama3-8B, Llama3-70B, Llama3-405B [35], Qwen2-7B, Qwen2-72B, Qwen1.5-110B [40], Phi3-small, Phi3-medium, Phi3.5-MoE [1], Mistral-7B, Mixtral-8x7B, and Mixtral-8x22B [17]. Unless specified otherwise, we assume a cluster size of 32 NPUs and we vary the batch size from 8 to 32 with increments of 8 and maximum sequence length from 2k to 16k with steps that correspond to the powers of 2k. In all experiments, we assume that both activations and weights are in int8 quantization.
Figure 5: Inference latency with respect to L2 cache size for various LLMs and their arithmeric mean, normalized to the smallest L2 cache size.
@@
Figure 6: Throughput density in terms of token/s/mm2 with respect to L2 cache size, with and without prefetching enabled. The curves have peak throughput density of 14.8 and 11.8 token/s/mm2 at 104 and 8 MB of L2 cache for with and without prefetching, respectively.
@@
Figure 7: Speedup obtained with prefetching over the baseline (left-side y-axis) as well as the allreduce and prefetch latencies for Attention and FFN layers (right-side y-axis) versus various number of (tensor-parallel) devices. The L2 cache for the accelerator is taken as 104 MB. The benchmark is a Llama3-405B model with a batch size of 16 and max. sequence length of 8k, where prompt and decode lengths are 2/3 and 1/3 of the total number of tokens, respectively.
@@
Figure 8: Speedup obtained with prefetching over the baseline for various LLMs with respect to the network bandwidth of the links between the devices. The L2 cache size and the number of devices are taken as 104 MB and 128, respectively. The batch size and max. sequence length are equal to 16 and 8k, where prompt and decode lengths are 2/3 and 1/3 of the total number of tokens, respectively.
Similar to the cluster size, the network bandwidth between the devices also has a considerable impact on the speedups achieved by the proposed method, as it directly affects the allreduce latencies. Figure 8 plots the speedups obtained by the proposed method for various LLMs with respect to the network bandwidth. We observe that the speedup generally increases with the network bandwidth for all models. This is due to the fact that increasing the bandwidth reduces the allreduce latencies, which enables to hide a larger portion of the communication overhead, leading to a higher speedup. On average, we observe that increasing the device-to-device link bandwidth from 200 to 1000 Gbps elevates the speedup from 1.18$\times$ to 1.27$\times$ for a cluster that consists of 128 devices.
In short, we perform a design space exploration using a performance model for AI accelerators to understand the impact of hardware design parameters and specifications on the performance and efficiency of the PRESERVE framework. We show that the optimal L2 cache size increases from 8 MB to 104 MB when prefetching is taken into account while designing the AI accelerators, which improves the throughput density by 1.25$\times$ even in spite of the increased silicon area due to larger L2 cache. We also investigated the sensitivity of the speedup of the proposed method to various hardware specifications and demonstrated that increasing the cube throughput, L2 bus bandwidth, and the interconnection network bandwidth helps to achieve higher speedups with the proposed method. We also concluded that the maximum speedup is achieved when the allreduce and prefetch latencies are equal to each other, which makes the speedup sensitive to the cluster size.
6 Related Work
KV-cache optimization & operator fusion
Due to the autoregressive nature of Transformer architectures, the performance of LLM inference is mostly bounded by off-chip memory bandwidth. To mitigate the memory bottlenecks, researchers have proposed several efficient kernel implementations. Flash Attention and Deepspeed-Inference resort to tiling and operator fusion to minimize the off-chip bandwidth usage and increase the efficiency of self-attention layers [7, 3]. Kwon et al. proposed PagedAttention, in which the memory management of KV-caches are optimized to improve the memory usage [19]. While these techniques may improve the single-device performance of Transformer architectures, in most cases, LLM inference still suffers from memory-bandwidth bottlenecks.
Distributed inference
Another major obstacle to LLM inference is the massive size of the models and KV-cache, which typically exceed the HBM capacity of a single accelerator for large models and long context sizes. To enable serving LLM inference for such large models with long context sizes, numerous prior work have focused on distributed LLM inference using a combination of parallelization techniques such as data-, tensor- [31], pipeline- [3, 26], expert- [9, 29, 32], and context-parallelism [22]. While these techniques might reduce the memory requirements and computational workload of LLM inference per device, the overhead of communication between the devices often places an upper limit on the scalability of distributed LLM inference systems.
To mitigate the communication overhead in distributed LLM inference, prior work have proposed fused kernels that overlap computation with communication [30, 39, 4]. In these methods, the compute and communication operations are partitioned into small chunks and processed in a pipeline fashion so that they overlap with each other. However, these fused kernels have several limitations. First, partitioning compute and communication operations into smaller chunks often reduces their computational efficiency due to factors such as short and irregular memory accesses. Second, this approach would require developing new kernels for each combination of compute and communication operation types, requiring significant engineering efforts and modifications in the user code. Third, the communication can be overlapped only with the immediate next compute operation, thus they do not allow overlapping multiple compute operations with the communication. For instance, these techniques can not overlap reading KV-cache with allreduce in an Attention layer because self-attention does not immediately follow allreduce in Transformer architectures. Therefore, these techniques do not provide a practical, effective, and complete solution towards hiding communication overhead in distributed LLM inference.
Other system-level optimizations
Besides kernel-level optimizations and parallelization techniques, several work have proposed system-level optimizations to improve the scalability and resource utilization of LLM inference systems. Several methods have been proposed to overlap compute-intensive prefill stage with memory-bound decode stage [41, 2]. Moreover, REEF introduces microsecond-scale kernel preemption to improve the overall throughput [10]. However, these methods address only memory bottlenecks and do not offer a solution to communication overheads. Alternatively, overlapping communication with compute by interleaving multiple requests has also been proposed [8]. Unfortunately, this approach requires splitting a batch of requests into two streams, which reduces the computational efficiency and resource contention in case compute and communication do not align perfectly.
To hide the memory and I/O latencies in LLM inference systems, software prefetching techniques have been used. PrefetchML proposes a software framework that prefetches large ML models from a database to the host memory [6]. ZeRO-Inference prefetches a number of layers ahead of time from CPU memory or NVMe to the GPU memory to hide the latency of communication between host and devices [3]. Similarly, DeepUM leverages correlation prefetching to hide the latency of page migration between GPU and CPU memories [18]. RaLMSpec introduces a speculative retrieval mechanism to predict and prefetch documents from a knowledge database to device memory in retrieval-augmented language models (RaLM) inference [42]. While these methods demonstrate the potential of software prefetching in LLM inference, none of these proposals leverage the on-chip L2 cache to maximize the data locality and do not address the challenges of limited L2 cache size for this purpose.
7 Conclusions
We performed a series of experiments on a commercial AI accelerator and showed that the proposed method achieves up to 1.6$\times$ end-to-end speedup on open-source state-of-the-art LLMs. Furthermore, we conducted a design space exploration to identify the optimal hardware design for AI accelerators with PRESERVE. We showed that, the optimal L2 cache size shifts from 8 MB to 104 MB when the proposed prefetching scheme is taken into consideration. Our analysis showed a further 1.25$\times$ improvement on performance per cost with PRESERVE for AI accelerators with the optimal L2 cache size.
Appendix
Figure 9: The average speedup obtained with prefetching over the baseline with respect to the cube throughput for various L2 bandwidth values. The L2 cache for the accelerator is taken as 104 MB and the cluster size is equal to 32. The batch size and max. sequence length are equal to 16 and 8k, where prompt and decode lengths are 2/3 and 1/3 of the total number of tokens, respectively.
Extending DSE to L2 bandwidth and compute units

### v2

@@ 5. Design Space Exploration
Table 2. Model parameters used in design space exploration.
@@
Table 2 summarizes the values of the model parameters used in our design space exploration. The silicon area for cores and L2 SRAM cache are taken from Lin et al. (Lin et al., 2020), whose design is fabricated in a 7nm technology node. We assumed that the accelerator has a theoretical peak throughput of 800 TeraOps/s, which is similar to existing inference accelerators (Huawei, 2022) (Li et al., 2024) and it is equipped with 4x HBM2e memory, which provides 64 GB of capacity and 1840 GB/s bandwidth in total (Synopsys, 2023).
For the design space exploration, we selected a number of open-source, state-of-the-art LLMs, namely Llama3-8B, Llama3-70B, Llama3-405B (Touvron et al., 2023), Qwen2-7B, Qwen2-72B, Qwen1.5-110B (Yang et al., 2024), Phi3-small, Phi3-medium, Phi3.5-MoE (Abdin et al., 2024), Mistral-7B, Mixtral-8x7B, and Mixtral-8x22B (Jiang et al., 2024). Unless specified otherwise, we assume a cluster size of 32 NPUs and we vary the batch size from 8 to 32 with increments of 8 and maximum sequence length from 2k to 16k with steps that correspond to the powers of 2k. In all experiments, we assume that both activations and weights are in int8 quantization.
Figure 6. Inference latency with respect to L2 cache size for various LLMs and their arithmeric mean, normalized to the smallest L2 cache size.
@@
Figure 7. Throughput density in terms of token/s/mm2 with respect to L2 cache size, with and without prefetching enabled. The curves have peak throughput density of 14.8 and 11.8 token/s/mm2 at 104 and 8 MB of L2 cache for with and without prefetching, respectively.
@@
Figure 8. Speedup obtained with prefetching over the baseline for various LLMs with respect to the network bandwidth of the links between the devices. The L2 cache size and the number of devices are taken as 104 MB and 128, respectively. The batch size and max. sequence length are equal to 16 and 8k, where prompt and decode lengths are 2/3 and 1/3 of the total number of tokens, respectively.
Besides the L2 cache, the network bandwidth between the devices also has a considerable impact on the speedups achieved by the proposed method, as it directly affects the allreduce latencies. Figure 8 plots the speedups obtained by the proposed method for various LLMs with respect to the network bandwidth. We observe that the speedup generally increases with the network bandwidth for all models. This is due to the fact that increasing the bandwidth reduces the allreduce latencies, which enables to hide a larger portion of the communication overhead, leading to a higher speedup. On average, we observe that increasing the device-to-device link bandwidth from 200 to 1000 Gbps elevates the speedup from 1.18$\times$ to 1.27$\times$ for a cluster that consists of 128 devices.
In short, we perform a design space exploration using a performance model for AI accelerators to understand the impact of hardware design parameters and specifications on the performance and efficiency of the PRESERVE framework. We show that the optimal L2 cache size increases from 8 MB to 104 MB when prefetching is taken into account while designing the AI accelerators, which improves the throughput density by 1.25$\times$ even in spite of the increased silicon area due to larger L2 cache.
6. Related Work
KV-cache optimization & operator fusion
Due to the autoregressive nature of Transformer architectures, the performance of LLM inference is mostly bounded by off-chip memory bandwidth. To mitigate the memory bottlenecks, researchers have proposed several efficient kernel implementations. Flash Attention and Deepspeed-Inference resort to tiling and operator fusion to minimize the off-chip bandwidth usage and increase the efficiency of self-attention layers (Dao et al., 2022; Aminabadi et al., 2022). Kwon et al. proposed PagedAttention, in which the memory management of KV-caches are optimized to improve the memory usage (Kwon et al., 2023). While these techniques may improve the single-device performance, in most cases, LLM inference still suffers from memory-bandwidth bottlenecks.
Distributed inference
Another major obstacle to LLM inference is the massive size of the models and KV-cache, which typically exceed the HBM capacity of a single accelerator for large models and long context sizes. To enable serving LLM inference for such large models with long context sizes, numerous prior work have focused on distributed LLM inference using a combination of parallelization techniques such as data-, tensor- (Shoeybi et al., 2019), pipeline- (Aminabadi et al., 2022; NVIDIA, 2023a), expert- (Fedus et al., 2022; Rajbhandari et al., 2022; Singh et al., 2023), and context-parallelism (Li et al., 2023). While these techniques might reduce the memory requirements and computational workload of LLM inference per device, the overhead of communication between the devices often places an upper limit on the scalability of distributed LLM inference systems.
To mitigate the communication overhead in distributed LLM inference, prior work have proposed fused kernels that overlap computation with communication (Rashidi et al., 2021; Hoefler and Lumsdaine, 2008; Punniyamurthy et al., 2024; Wang et al., 2023; Chang et al., 2024). In these methods, the compute and communication operations are partitioned into small chunks and processed in a pipeline fashion so that they overlap with each other. However, these fused kernels have several limitations. First, partitioning compute and communication operations into smaller chunks often reduces their computational efficiency due to factors such as short and irregular memory accesses. Second, this approach would require developing new kernels for each combination of compute and communication operation types, requiring significant engineering efforts. Third, the communication can be overlapped only with the immediate next compute operation, thus they do not allow overlapping multiple compute operations with the communication. For instance, these techniques can not overlap reading KV-cache with allreduce in an Attention layer because self-attention does not immediately follow allreduce in Transformer architectures. Therefore, these techniques do not provide a practical, effective, and complete solution towards hiding communication overhead in distributed LLM inference.
Other system-level optimizations
Besides kernel-level optimizations and parallelization techniques, several work have proposed system-level optimizations to improve the scalability and resource utilization of LLM inference systems. Varios methods have been proposed to overlap compute-intensive prefill stage with memory-bound decode stage (Yu et al., 2022; Agrawal et al., 2023). Moreover, REEF introduces microsecond-scale kernel preemption to improve the overall throughput (Han et al., 2022). However, these methods address only memory bottlenecks and do not offer a solution to communication overheads. Alternatively, overlapping communication with compute by interleaving multiple requests has also been proposed (Du et al., 2024). Unfortunately, this approach requires splitting a batch of requests into two streams, which leads to the computational inefficiencies and resource contention in case compute and communication do not align perfectly.
To hide the memory and I/O latencies in LLM inference systems, software prefetching techniques have been used. PrefetchML proposes a software framework that prefetches large ML models from a database to the host memory (Daniel et al., 2016). ZeRO-Inference prefetches a number of layers ahead of time from CPU memory or NVMe to the GPU memory to hide the latency of communication between host and devices (Aminabadi et al., 2022). Similarly, DeepUM leverages correlation prefetching to hide the latency of page migration between GPU and CPU memories (Jung et al., 2023). RaLMSpec introduces a speculative retrieval mechanism to predict and prefetch documents from a knowledge database to device memory in retrieval-augmented language models (RaLM) inference (Zhang et al., 2024). While these methods demonstrate the potential of software prefetching in LLM inference, none of these proposals leverage the on-chip L2 cache to maximize the data locality and do not address the challenges of limited L2 cache size for this purpose.
7. Limitations of the Proposed Method
While the previous sections have demonstrated the potential benefits, we now discuss the known limitations of the proposed method. First, the effectiveness of the proposed method is limited by how much model weights and KV cache can fit in the on-chip memory capacity of the AI accelerators. Therefore, any factor that increases the model weight size and KV-cache per layer and per device (e.g., batch size, sequence length, precision) may affect the speedup obtained with the proposed method. However, tight latency constraints in online inference applications (e.g., chatbots, code assistants) often place an upper bound on the batch size. Moreover, for long sequence lengths, sequence parallelism is often used, which reduces KV-cache size per device. Finally, more aggressive and efficient quantization methods (Lin et al., 2024; Hooper et al., 2024) are being developed, which further reduces the memory requirements with minimal impact on model accuracy. Therefore, the proposed method remains effective and applicable to a wide range of application settings.
8. Conclusions
We performed a series of experiments on a commercial AI accelerator and showed that the proposed method achieves up to 1.6$\times$ end-to-end speedup on state-of-the-art LLMs. Furthermore, we conducted a design space exploration to identify the optimal hardware design for AI accelerators with PRESERVE. We showed that, the optimal L2 cache size shifts from 8 MB to 104 MB when the proposed prefetching scheme is taken into consideration. Our analysis showed a further 1.25$\times$ improvement on performance per cost with PRESERVE for AI accelerators with the optimal L2 cache size.
Appendix
Figure 9. The average speedup obtained with prefetching over the baseline with respect to the cube throughput for various L2 bandwidth values. The L2 cache for the accelerator is taken as 104 MB and the cluster size is equal to 32. The batch size and max. sequence length are equal to 16 and 8k, where prompt and decode lengths are 2/3 and 1/3 of the total number of tokens, respectively.
Extending DSE to L2 bandwidth and compute units
@@ 8. Conclusions
Figure 10. Speedup obtained with prefetching over the baseline (left-side y-axis) as well as the allreduce and prefetch latencies for Attention and FFN layers (right-side y-axis) versus various number of (tensor-parallel) devices. The L2 cache for the accelerator is taken as 104 MB. The benchmark is a Llama3-405B model with a batch size of 16 and max. sequence length of 8k, where prompt and decode lengths are 2/3 and 1/3 of the total number of tokens, respectively.